3. Solve for y: 3(5y-2)-(y+4) = 24
Answer by checkley77(12844)

3(5y-2)-(y+4) = 24
15y-6-y-4=24
14y=24+6+4
14y=34
y=34/14
y=17/7 ans.
Proof:
3(5*17/7-2)-(17/7+4)=24
3(85/7-14/7)-(17/7+28/7)=24
3(71/7)-(45/7)=24
213/7-45/7=24
168/7=24
24=24